# On absolutely continuous spectrum for one-channel unitary operators

### v1

@@ Abstract.
Acknowledgement: This work has been supported by the Chilean grants FONDECYT 1230949, 1211576, 1211189, 1201836.
@@ 1. Setup and result
In recent years there has been some interest in the use of ‘reduced transfer matrices’ for Schrödinger operators, where the hopping between ‘shells’ or ‘slices’ of some graph are of fixed (but not full) rank [19, 47, 48, 50]. These matrices were introduced independently by Dwivedi-Chua [19] and Sadel [47, 48, 49]. In the strip case, ‘reduced’ means for instance that the dimension of the transfer matrices is smaller than the usual ‘twice of the strip width’, it is only ‘twice the rank’. In the most extreme case, where the connections are only of rank 1, we have $2\times 2$ transfer matrices. Such Hermitian operators were called ‘one-channel operators’ in [48]. This can occur even if the size of the ‘shells’ grow, leading to graphs with a radial growth corresponding to some higher dimension. The size reduction of the transfer matrices can help to analyze certain spectral aspects more easily, see for instance [47, 48].
@@ 1.1. One-channel unitary operators
Mostly, the direct some has to be understood as an Hilbert-space orthogonal sum. However, the operators we consider are of finite hopping type and extend naturally to the set of all functions from ${\mathbb{G}}$ to ${\mathbb{C}}$, and we may also use the notation above for $\Psi\in{\mathbb{C}}^{\mathbb{G}}$ with $\Psi_{n}=\Psi|_{{\mathbb{S}}_{n}}$ being the restriction of $\Psi$ to ${\mathbb{S}}_{n}$. In physics literature the space ${\mathbb{C}}^{\mathbb{G}}$ maybe referred as ‘generalized states’ and solutions to $U\Psi=z\Psi$ for $\Psi\in{\mathbb{C}}^{\mathbb{G}}$ as ‘generalized eigenfunctions’. Furthermore, for $n\leq N$, an element $\varphi\in{\mathbb{C}}^{{\mathbb{S}}_{n}}$ can also be considered as an element of ${\mathbb{C}}^{{\mathbb{G}}_{N}}$ or $\ell^{2}({\mathbb{G}})$, identifying $\varphi$ with $\varphi\oplus\bigoplus_{k\neq n}\mathbf{0}$. Using some adequate basis we will identify ${\mathbb{C}}^{{\mathbb{S}}_{n}}$ with ${\mathbb{C}}^{|{\mathbb{S}}_{n}|}$ later on, but we prefer the notation ${\mathbb{C}}^{{\mathbb{S}}_{n}}$ to distinguish the spaces for different shells with possibly same number of elements. Similarly, by notations like ${\mathbb{C}}^{{\mathbb{S}}_{n}\times l}$ we understand the set of linear maps from ${\mathbb{C}}^{l}$ to ${\mathbb{C}}^{S_{n}}$, which (given a basis of ${\mathbb{C}}^{{\mathbb{S}}_{n}}$) can be identified with the set of $|{\mathbb{S}}_{n}|\times l$ matrices, or with the set of $l$-tuples of vectors in ${\mathbb{C}}^{{\mathbb{S}}_{n}}$. First, we define the operator ${\mathcal{V}}$ by
@@
where ${\mathcal{P}}_{{\mathbb{S}}_{n}}$ is the orthogonal projection of $\ell^{2}({\mathbb{G}})$ onto ${\mathbb{C}}^{{\mathbb{S}}_{n}}=\ell^{2}({\mathbb{S}}_{n})$. Here, we use the standard physics scalar product $\langle\varphi,\varphi^{\prime}\rangle=\sum_{x}\bar{\varphi}(x)\varphi^{\prime}(x)$ and clearly, ${\mathcal{V}}$ is unitary.
@@
${\mathcal{W}}^{(u)}\,=\,u\,e_{(n,-)}e_{(n,-)}^{*}+P_{0}+\sum_{n=1}^{\infty}\left((e_{(n-1,+)},\;e_{(n,-)})W_{n}\begin{pmatrix}e_{(n-1,+)}^{*}\\ e_{(n,-)}^{*}\end{pmatrix}+P_{n}\right),$ (1.2)
where $u\in{\mathbb{U}}(1)=\partial{\mathbb{D}}$ is some sort of ’left boundary condition’, and $W_{n}\in{\mathbb{U}}(2)$. Here, ${\mathbb{U}}({\mathbb{S}}_{n})$ denotes the unitary operators on $\ell^{2}({\mathbb{S}}_{n})$ and ${\mathbb{U}}(k)$ the unitary operators on ${\mathbb{C}}^{k}$. In the notation for ${\mathcal{W}}$ above we interpret the vectors $e_{(n,\pm)}$ in ${\mathbb{C}}^{{\mathbb{S}}_{n}}$ as column vectors in $\ell^{2}({\mathbb{G}})$ and in the sense of matrices as maps from ${\mathbb{C}}$ to ${\mathbb{C}}^{{\mathbb{S}}_{n}}\subset\ell^{2}({\mathbb{G}})$, and $e_{(n,\pm)}^{*}$ as maps from ${\mathbb{C}}^{{\mathbb{S}}_{n}}$ or $\ell^{2}({\mathbb{G}})$ to ${\mathbb{C}}$. Also, $P_{n}$ is naturally interpreted as an operator on $\ell^{2}({\mathbb{G}})$ identifying it with $P_{n}\oplus\bigoplus_{k\neq n}\mathbf{0}$. Another way of representing ${\mathcal{W}}$ is by using the projections
@@
${\mathcal{W}}\,=\,{\mathcal{Q}}^{\perp}\,\oplus\,u\,\oplus\,\bigoplus_{n\in{\mathbb{Z}}_{+}^{*}}W_{n},$
@@
$W_{n}\,=\,{\mathcal{Q}}_{n}{\mathcal{W}}{\mathcal{Q}}_{n}\,\in\,{\mathbb{U}}(2)\quad\text{or $n\geq 1$ and}\quad u={\mathcal{Q}}_{0}{\mathcal{W}}{\mathcal{Q}}_{0}\,\in\,{\mathbb{U}}(1)\,.$
@@
Using an ortonormal basis of ${\mathbb{C}}^{{\mathbb{S}}_{n}}=\ell^{2}({\mathbb{S}}_{n})$ where $e_{(n,-)}$ is the first, and $e_{(n,+)}$ the last vector, one may write
@@
Basically, the sequence $W_{n}$ connecting the vectors $e_{(n,+)},e_{(n+1,-)}$ defines a ’channel’ through which waves can travel towards infinity (across the shells) when, considering the dynamics ${\mathcal{U}}^{k},\tilde{\mathcal{U}}^{k}$, $k\in{\mathbb{N}}$, $k\to\infty$.
@@
One may write a one-dimensional random quantum walk in this framework as the operator ${\mathcal{U}}$ using shells of size $|{\mathbb{S}}_{n}|=2$, (so $P_{n}=\mathbf{0}$) where $W_{n}$ are transpositions, $W_{n}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$. The definitions of the transfer matrices are guided by considering this case. In fact, one may think of ${\mathcal{U}}$ as a one-channel quantum walk and $V_{n}$ as ‘coin’ matrices in ${\mathbb{S}}_{n}$. Then the ‘walk’ happens according to the matrices $W_{n}$ across the channel $(e_{(n,+)},e_{(n+1,-)})_{n\in{\mathbb{N}}_{0}}$. Furthermore, in the case $|{\mathbb{S}}_{n}|=2$ the operator $\tilde{\mathcal{U}}$ corresponds exactly to the scattering zippers with size $L=2$ in [42]. As such scattering zippers include the case of CMV matrices and one-dimensional quantum walks, our framework can be seen as an extension of both.
Comments.
@@ 1.2. Transfer matrices
${\mathbb{U}}(1,1)\,=\,\left\{T\in{\mathbb{C}}^{2\times 2}\,:\,T^{*}\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}T=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}\right\}:$
@@
    ${\mathcal{U}}^{(u)}\Psi=z\Psi\;\wedge\;{\mathcal{W}}^{(u)}\Phi\,=\,\Psi$.
@@
    ${\mathcal{V}}\Psi=z\Phi\;\wedge\;{\mathcal{W}}^{(u)}\Phi\,=\,\Psi$.
@@
    $\tilde{\mathcal{U}}^{(u)}\Phi=z\Phi\,\wedge\;{\mathcal{W}}^{(u)}\Phi\,=\,\Psi$.
@@
$\begin{pmatrix}\Psi_{(n,-)}\\ \Psi_{(n,+)}\end{pmatrix}\,=\,\begin{pmatrix}\alpha_{z,n}&\beta_{z,n}\\ \gamma_{z,n}&\delta_{z,n}\end{pmatrix}\begin{pmatrix}\Phi_{(n,-)}\\ \Phi_{(n,+)}\end{pmatrix}$ (1.5)
@@
in case that $z^{-1}V_{n}-P_{n}$ is invertible (which is the case for all $10<|z|<$ 1). We also note that for $|z|=1$ this matrix is unitary by Proposition A.3. Clearly, it is a meromorphic function in $z$ by Cramer’s rule and using Proposition A.3, none of the poles lies on the unit circle. Moreover it is easy to see that in the limit $z\to 0$ one obtains the zero matrix. Hence, after analytic continuation, it is well defined for all $|z|\leq 1$.
@@
This is well defined for all $|z|>1$. In case where all inverses exist, one thus gets
@@
which is a special case of Proposition A.3 and which shows unitarity in the case $|z|=1$ in another way. The guide for defining the transfer matrices is the special case where all $W_{n}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ as it will be the case for a 1D quantum walk, cf. subsection 2.1. In this case $\Psi_{(n+1,-)}=\Phi_{(n,+)}$, $\Psi_{(n-1,+)}=\Phi_{(n,-)}$, and it makes sense to define a transfer matrix associated to $V_{n}$ by
$\begin{pmatrix}\Phi_{(n,+)}\\ \Psi_{(n,+)}\end{pmatrix}\,=\,T_{z,n}^{\sharp}\,\begin{pmatrix}\Psi_{(n,-)}\\ \Phi_{(n,-})\end{pmatrix},$ (1.8)
where by (1.5) and Proposition 1.1 $T^{\sharp}_{z,n}$ exists if $\beta_{z,n}\neq 0$ (which is equivalent to $\tilde{\beta}_{z,n}\neq 0$), in which case
@@
In order to complete to a transfer matrix coming from the ’level before’, we define the transfer matrix associated to $W_{n}$ by
@@
The equation $\Psi_{(0,-)}=u\Phi_{(0,-)}$ for the operator pair $({\mathcal{U}}_{u},\tilde{\mathcal{U}}_{u})$ can be understood as some boundary condition. Here, we will not incorporate the boundary condition into the transfer matrices and simply define
@@
Thus, in this case, for any $n$, $T^{\sharp}_{z,n}$ is well defined except for finitely many $z$, as $\beta_{z,n}$ is meromorphic.
@@
and the previous statement gives $e_{(n,+)}^{*}V_{n}^{m+1}e_{(n,-)}=0$ if and only if $e_{(n,+)}^{*}V_{n}(P_{n}V_{n})^{m}e_{(n,-)}$. which finishes the induction step as $e_{(n,-)}^{*}V_{n}^{l}e_{(n,+)}=0$ for $l=0,\ldots,m$ by induction hipothesis. ∎
On the set where $Q_{n}^{*}(z^{-1}V_{n}-P_{n})^{-1}Q_{n}$ is not defined we may use analytic continuation in $z$ to define $T^{\sharp}_{z,n}$. By the comments above, one can use the formula (1.9), for $0<|z|\leq 1$ where $\beta_{z,n}\neq 0$, and one can use (1.10) for $|z|>1$ where $\tilde{\beta}_{z,n}\neq 0$. The exceptional set where $T^{\sharp}_{z,n}$ and $T_{z,n}$ are not defined is thus given by
@@
Let $\mu^{(u)}$ denote the spectral measure of the operator ${\mathcal{U}}^{(u)}$ (o alternatively $\tilde{\mathcal{U}}^{(u)}$) at the vector $e_{(0,-)}$, meaning
@@
The second equation follows easily as $\tilde{\mathcal{U}}^{(u)}=({\mathcal{W}}^{(u)})^{*}{\mathcal{U}}^{(u)}\,{\mathcal{W}}^{(u)}$ and ${\mathcal{W}}^{(u)}e_{(0,-)}=u\,e_{(0,-)}$ with $|u|=1$. We get some analogue formula as in the Hermitian case [48, Theorem 2]. Similar, we need to separate the part induced by compactly supported eigenfunctions. First we define
@@
Let us note that for some particular eigenvalue $z_{0}\in\partial{\mathbb{D}}$, the eigenfunction in the intersection of the cyclic space of $e_{(0,-)}$ with ${\mathbb{H}}_{c}^{(u)}$ giving $\nu^{(u)}(f)$ may not be compactly supported, but in such a case, it must be the limit of compactly supported eigenfunctions of the same eigenvalue. This only may happen if $z_{0}$ is an eigenvalue of infinite multiplicity.
@@
For any $u\in{\mathbb{U}}(1)$ there is a point measure $\nu^{(u)}$ supported on ${\mathcal{A}}$ such that for $f\in C(\partial{\mathbb{D}})=C({\mathbb{U}}(1))$ we have
@@
Assume that for $p>2$, and $\varphi_{0}<\varphi_{1}$ one has
@@
then, for any $u\in{\mathbb{U}}(1)$, the positive measure $\mu^{(u)}-\nu^{(u)}$ is purely absolutely continuous in $e^{i(\varphi_{0},\varphi_{1})}=\{e^{i\varphi}\,:\,\varphi\in(\varphi_{0},\varphi_{1})\}$ w.r.t. the Haar measure on $\partial{\mathbb{D}}={\mathbb{U}}(1)$ and
@@
(C1) The pairs $(\widehat{V}_{n},\widehat{W}_{n})_{n=0}^{\infty}$ are independent.
@@
Assume that (C1), (C2), (C3) hold. Then, there is a set $\Omega^{\prime}$ of probability one, ${\mathbb{P}}(\Omega^{\prime})=1$ such that for all $\omega\in\Omega^{\prime}$, the spectrum of ${\mathcal{U}}_{\omega}$ is purely absolutely continuous in $\Sigma$, and, $\sigma_{ac}({\mathcal{U}}_{\omega})\,=\,\overline{\Sigma}$.
Comments.
@@
    for any $z\in\Sigma$, uniformly for $z$ in compact subsets of $\Sigma$. Assumptions (C1), (C2), (C3) guarantee this. Without assumption (C3) one could have $\hat{b}_{n}$ or $\hat{\mathfrak{b}}_{n}$ closer and closer to cero with smaller and smaller positive probabilities, such that (C2) is satisfied but not (C3) and also not the technical condition needed as stated above, as the inverses of $\hat{b}_{n},\hat{\mathfrak{b}}_{n}$ appear in the transfer matrices.

### v2

@@ Abstract.
corresponding author: C. Sadel.
@@ 1. Setup and result
In recent years there has been some interest in the use of ‘reduced transfer matrices’ for Schrödinger operators, where the hopping between ‘shells’ or ‘slices’ of some graph are of fixed (but not full) rank [19, 47, 48, 50]. These matrices were introduced independently by Dwivedi-Chua [19] and Sadel [47, 48, 49]. In the strip case, ‘reduced’ means for instance that the dimension of the transfer matrices is smaller than the usual ‘twice of the strip width’, it is only ‘twice the rank’. In the most extreme case, where the connections are only of rank 1, we have $2\times 2$ transfer matrices. Such Hermitian operators were called ‘one-channel operators’ in [48]. This can occur even if the size of the ‘shells’ grow, leading to graphs with a radial growth corresponding to some higher dimensions. The size reduction of the transfer matrices can help to analyze certain spectral aspects more easily, see for instance [47, 48].
@@ 1.1. One-channel unitary operators
Mostly, the direct sum has to be understood as an Hilbert-space orthogonal sum. However, the operators we consider are of finite hopping type and extend naturally to the set of all functions from ${\mathbb{G}}$ to ${\mathbb{C}}$, and we may also use the notation above for $\Psi\in{\mathbb{C}}^{\mathbb{G}}$ with $\Psi_{n}=\Psi|_{{\mathbb{S}}_{n}}$ being the restriction of $\Psi$ to ${\mathbb{S}}_{n}$. In physics literature the space ${\mathbb{C}}^{\mathbb{G}}$ maybe referred as ‘generalized states’ and solutions to $U\Psi=z\Psi$ for $\Psi\in{\mathbb{C}}^{\mathbb{G}}$ as ‘generalized eigenfunctions’. Furthermore, for $n\leq N$, an element $\varphi\in{\mathbb{C}}^{{\mathbb{S}}_{n}}$ can also be considered as an element of ${\mathbb{C}}^{{\mathbb{G}}_{N}}$ or $\ell^{2}({\mathbb{G}})$, identifying $\varphi$ with $\varphi\oplus\bigoplus_{k\neq n}\mathbf{0}$. Using some adequate basis we will identify ${\mathbb{C}}^{{\mathbb{S}}_{n}}$ with ${\mathbb{C}}^{|{\mathbb{S}}_{n}|}$ later on, but we prefer the notation ${\mathbb{C}}^{{\mathbb{S}}_{n}}$ to distinguish the spaces for different shells with possibly same number of elements. Similarly, by notations like ${\mathbb{C}}^{{\mathbb{S}}_{n}\times l}$ we understand the set of linear maps from ${\mathbb{C}}^{l}$ to ${\mathbb{C}}^{S_{n}}$, which (given a basis of ${\mathbb{C}}^{{\mathbb{S}}_{n}}$) can be identified with the set of $|{\mathbb{S}}_{n}|\times l$ matrices, or with the set of $l$-tuples of vectors in ${\mathbb{C}}^{{\mathbb{S}}_{n}}$. First, we define the operator ${\mathcal{V}}$ by
@@
where ${\mathcal{P}}_{{\mathbb{S}}_{n}}$ is the orthogonal projection of $\ell^{2}({\mathbb{G}})$ onto ${\mathbb{C}}^{{\mathbb{S}}_{n}}=\ell^{2}({\mathbb{S}}_{n})$. Here, we use the standard physics scalar product $\langle\varphi,\varphi^{\prime}\rangle=\sum_{x}\bar{\varphi}(x)\varphi^{\prime}(x)$, and, moreover, ${\mathbb{U}}({\mathbb{S}}_{n})$ denotes the unitary matrices on ${\mathbb{C}}^{{\mathbb{S}}_{n}}$. Thus, ${\mathcal{V}}$ is unitary.
@@
${\mathcal{W}}^{(u)}\,=\,u\,e_{(n,-)}e_{(n,-)}^{*}+P_{0}+\sum_{n=1}^{\infty}\left((e_{(n-1,+)},\;e_{(n,-)})W_{n}\begin{pmatrix}e_{(n-1,+)}^{*}\\ e_{(n,-)}^{*}\end{pmatrix}+P_{n}\right).$ (1.2)
Here, $u\in{\mathbb{U}}(1)=\partial{\mathbb{D}}$ is some sort of ‘left boundary condition’, and $W_{n}\in{\mathbb{U}}(2)$, where ${\mathbb{U}}(k)$ denotes the unitary operators on ${\mathbb{C}}^{k}$. In the notation for ${\mathcal{W}}$ above we interpret the vectors $e_{(n,\pm)}$ in ${\mathbb{C}}^{{\mathbb{S}}_{n}}$ as column vectors in $\ell^{2}({\mathbb{G}})$ and in the sense of matrices as maps from ${\mathbb{C}}$ to ${\mathbb{C}}^{{\mathbb{S}}_{n}}\subset\ell^{2}({\mathbb{G}})$, and $e_{(n,\pm)}^{*}$ as maps from ${\mathbb{C}}^{{\mathbb{S}}_{n}}$ or $\ell^{2}({\mathbb{G}})$ to ${\mathbb{C}}$. Also, $P_{n}$ is naturally interpreted as an operator on $\ell^{2}({\mathbb{G}})$ identifying it with $P_{n}\oplus\bigoplus_{k\neq n}\mathbf{0}$. Another way of representing ${\mathcal{W}}$ is by using the projections
@@
${\mathcal{W}}^{(u)}\,=\,{\mathcal{Q}}^{\perp}\,\oplus\,u\,\oplus\,\bigoplus_{n\in{\mathbb{Z}}_{+}^{*}}W_{n},$
@@
$W_{n}\,=\,{\mathcal{Q}}_{n}{\mathcal{W}}^{(u)}{\mathcal{Q}}_{n}\,\in\,{\mathbb{U}}(2)\quad\text{or $n\geq 1$ and}\quad u={\mathcal{Q}}_{0}{\mathcal{W}}^{(u)}{\mathcal{Q}}_{0}\,\in\,{\mathbb{U}}(1)\,.$
@@
Using an orthonormal basis of ${\mathbb{C}}^{{\mathbb{S}}_{n}}=\ell^{2}({\mathbb{S}}_{n})$ where $e_{(n,-)}$ is the first, and $e_{(n,+)}$ the last vector, one may write
@@
Basically, the sequence $W_{n}$ connecting the vectors $e_{(n,+)},e_{(n+1,-)}$ defines a ‘channel’ through which waves can travel towards infinity (across the shells) when considering the dynamics ${\mathcal{U}}^{k},\tilde{\mathcal{U}}^{k}$, $k\in{\mathbb{N}}$, $k\to\infty$.
@@
One may write a one-dimensional random quantum walk in this framework as the operator ${\mathcal{U}}$ using shells of size $|{\mathbb{S}}_{n}|=2$, (so $P_{n}=\mathbf{0}$) where $W_{n}$ are transpositions, $W_{n}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$. The definitions of the transfer matrices are guided by considering this case. In fact, one may think of ${\mathcal{U}}$ as a one-channel quantum walk and $V_{n}$ as ‘coin’ matrices in ${\mathbb{S}}_{n}$. Then the ‘walk’ happens according to the matrices $W_{n}$ across the channel $(e_{(n,+)},e_{(n+1,-)})_{n\in{\mathbb{N}}_{0}}$. Furthermore, in the case $|{\mathbb{S}}_{n}|=2$ the operator $\tilde{\mathcal{U}}$ corresponds exactly to the scattering zippers with size $L=2$ in [42], which include CMV matrices and one-dimensional quantum walks. Thus, our framework can be seen as an extension of both.
Remark 1.1.
@@ 1.2. Transfer matrices
${\mathbb{U}}(1,1)\,=\,\left\{T\in{\mathbb{C}}^{2\times 2}\,:\,T^{*}\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}T=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}\right\}.$
@@
    ${\mathcal{U}}^{(u)}\Psi=z\Psi\quad\text{{\rm and}}\quad{\mathcal{W}}^{(u)}\Phi\,=\,\Psi$.
@@
    ${\mathcal{V}}\Psi=z\Phi\quad\text{{\rm and}}\quad{\mathcal{W}}^{(u)}\Phi\,=\,\Psi$.
@@
    $\tilde{\mathcal{U}}^{(u)}\Phi=z\Phi\,\quad\text{{\rm and}}\quad{\mathcal{W}}^{(u)}\Phi\,=\,\Psi$.
@@
$\begin{pmatrix}\Psi_{(n,-)}\\ \Psi_{(n,+)}\end{pmatrix}\,=\,\begin{pmatrix}\alpha_{z,n}&\beta_{z,n}\\ \gamma_{z,n}&\delta_{z,n}\end{pmatrix}\begin{pmatrix}\Phi_{(n,-)}\\ \Phi_{(n,+)}\end{pmatrix},$ (1.5)
@@
in case that $z^{-1}V_{n}-P_{n}$ is invertible. We also note that for $|z|=1$ the matrix defined in (1.2) is unitary by part b) of Proposition A.3, where
$A=Q_{n}^{*}z^{-1}V_{n}Q_{n},\,B=Q_{n}^{*}z^{-1}V_{n}Q_{n}^{\perp},\,C=(Q_{n}^{\perp})^{*}z^{-1}V_{n}Q_{n},\,D=(Q_{n}^{\perp})^{*}z^{-1}V_{n}Q_{n}^{\perp}\,,$
$P=(Q_{n}^{\perp})^{*}P_{n}Q_{n}^{\perp}\,=\,(Q_{n}^{\perp})^{*}Q_{n}^{\perp}\,=\,\mathbf{I}\,.$
Here, the column vectors of $Q_{n}^{\perp}\in{\mathbb{C}}^{{\mathbb{S}}_{n}\times(|{\mathbb{S}}_{n}|-2)}$ complete the columns of $Q_{n}$ to an orthonormal basis of ${\mathbb{C}}^{{\mathbb{S}}_{n}}$.
Remark 1.4.
Clearly, $z\mapsto Q_{n}^{*}(z^{-1}V_{n}-P_{n})^{-1}Q_{n}$ is a rational function by Cramer’s rule, it exists for all $10<|z|<$ 1. Using Proposition A.3, none of the poles lies on the unit circle. Rewriting $(z^{-1}V_{n}-P_{n})^{-1}=z(V_{n}-zP_{n})^{-1}$, it is easy to see that in the limit $z\to 0$ one obtains the zero matrix. Hence, after analytic continuation, $\alpha_{z,n},\,\beta_{z,n},\,\gamma_{z,n}$ and $\delta_{z,n}$ are well defined for all $|z|\leq 1$.
@@
This is well defined for all $|z|>1$. In the case where all inverses exist, one thus gets
@@
which is a special case of Proposition A.3 c) and also shows unitarity for $|z|=1$. The guide for defining the transfer matrices is the special case where all $W_{n}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ as it will be the case for a 1D quantum walk, cf. subsection 2.1.
For this choice of $W_{n}$ one obtains $\Psi_{(n+1,-)}=\Phi_{(n,+)}$, $\Psi_{(n-1,+)}=\Phi_{(n,-)}$, and it makes sense to define a transfer matrix associated to $V_{n}$ by
$\begin{pmatrix}\Phi_{(n,+)}\\ \Psi_{(n,+)}\end{pmatrix}\,=\,T_{z,n}^{\sharp}\,\begin{pmatrix}\Psi_{(n,-)}\\ \Phi_{(n,-})\end{pmatrix}.$ (1.8)
By (1.5) and Proposition 1.2 $T^{\sharp}_{z,n}$ exists if $\beta_{z,n}\neq 0$ (or $\tilde{\beta}_{z,n}\neq 0$), in which case
@@
In order to complete to a transfer matrix coming from the ‘level before’, we define the transfer matrix associated to $W_{n}$ by
@@
Note that for the special choice $W_{n}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ we have $T_{n}^{\flat}=\mathbf{I}_{2}$.
The equation $\Psi_{(0,-)}=u\Phi_{(0,-)}$ for the operator pair $({\mathcal{U}}^{(u)},\tilde{\mathcal{U}}^{(u)})$ can be understood as some boundary condition. Here, we will not incorporate the boundary condition into the transfer matrices and simply define
@@
Thus, in this case, for any $n$, $T^{\sharp}_{z,n}$ is well defined except for finitely many $z$, as $z\mapsto\beta_{z,n}$ is a rational function.
@@
and the previous statement gives $e_{(n,+)}^{*}V_{n}^{m+1}e_{(n,-)}=0$ if and only if $e_{(n,+)}^{*}V_{n}(P_{n}V_{n})^{m}e_{(n,-)}=0$. which finishes the induction step as $e_{(n,-)}^{*}V_{n}^{l}e_{(n,+)}=0$ for $l=0,\ldots,m$ by induction hipothesis. ∎
On the set where either $(z^{-1}V_{n}-P_{n})^{-1}$ or $\varphi_{\sharp}\big{(}Q_{n}^{*}(z^{-1}V_{n}-P_{n})^{-1}Q_{n}\big{)}$ is not defined we may use analytic continuation in $z$ to define $T^{\sharp}_{z,n}$ wherever possible. By Remark 1.4 one can use the formula (1.9) for $0<|z|\leq 1$ where $\beta_{z,n}\neq 0$, and one can use (1.10) for $|z|>1$ where $\tilde{\beta}_{z,n}\neq 0$. The exceptional set where $T^{\sharp}_{z,n}$ and $T_{z,n}$ are not defined is thus given by
@@
Let $\mu^{(u)}$ denote the spectral measure of the operator ${\mathcal{U}}^{(u)}$ (or alternatively $\tilde{\mathcal{U}}^{(u)}$) at the vector $e_{(0,-)}$, meaning
@@
The second equation follows easily as $\tilde{\mathcal{U}}^{(u)}=({\mathcal{W}}^{(u)})^{*}{\mathcal{U}}^{(u)}\,{\mathcal{W}}^{(u)}$ and ${\mathcal{W}}^{(u)}e_{(0,-)}=u\,e_{(0,-)}$ with $|u|=1$.
Similarly to the Hermitian case (cf. [48]) we need to separate the measure part induced by compactly supported eigenfunctions. First we define
@@
Remark 1.7.
Note that for some particular eigenvalue $z_{0}\in{\mathbb{U}}(1)$, the eigenfunction in the intersection of the cyclic space of $e_{(0,-)}$ with ${\mathbb{H}}_{c}^{(u)}$ giving $\nu^{(u)}(f)$ may not be compactly supported, but in such a case, it must be the limit of compactly supported eigenfunctions of the same eigenvalue. This only may happen if $z_{0}$ is an eigenvalue of infinite multiplicity.
We obtain the following analogue to [48, Theorem 2].
@@
For any $u\in{\mathbb{U}}(1)$, $\nu^{(u)}$ is supported on ${\mathcal{A}}$ and for $f\in C(\partial{\mathbb{D}})=C({\mathbb{U}}(1))$ we have
@@
Assume that for $p>1$, and $\varphi_{0}<\varphi_{1}$ one has
@@
then, for any $u\in{\mathbb{U}}(1)$, the positive measure $\mu^{(u)}-\nu^{(u)}$ is purely absolutely continuous in $e^{i(\varphi_{0},\varphi_{1})}=\{e^{i\varphi}\,:\,\varphi\in(\varphi_{0},\varphi_{1})\}$ w.r.t. the Haar measure on $\partial{\mathbb{D}}={\mathbb{U}}(1)$, has density in $L^{p}(e^{i(\varphi_{0},\varphi_{1})})$, and
@@
(C1) The family of pairs $\{(\widehat{V}_{n},\widehat{W}_{n})\}_{n=0}^{\infty}$ is independent222equivalently, one may state that the family of transfer matrices $(\widehat{T}_{z,n})_{n}$ is independent, $\widehat{V}_{n}$ and $\widehat{W}_{n}$ may have correlations. .
@@
Assume that (C1), (C2), (C3) hold. Then, there is a set $\Omega^{\prime}$ of probability one, ${\mathbb{P}}(\Omega^{\prime})=1$, such that for all $\omega\in\Omega^{\prime}$, the spectrum of ${\mathcal{U}}_{\omega}$ is purely absolutely continuous in $\Sigma$, and, $\sigma_{ac}({\mathcal{U}}_{\omega})\,=\,\overline{\Sigma}=\sigma_{ess}({\mathcal{U}}_{\omega})$.
Remark 1.9.
@@
    for any $z\in\Sigma$, uniformly for $z$ in compact subsets of $\Sigma$. Assumptions (C1), (C2), (C3) guarantee this. Without assumption (C3) one could have $\hat{b}_{n}$ or $\hat{\mathfrak{b}}_{n}$ closer and closer to zero with smaller and smaller positive probabilities, such that (C2) is satisfied but not (C3) and also not the technical condition needed as stated above, as the inverses of $\hat{b}_{n},\hat{\mathfrak{b}}_{n}$ appear in the transfer matrices.